OpenStudy (theedwardsfamily):

Will medal and fan!!!!!! :DDDDD Which equations model this problem? A recipe that makes 1 serving calls for 0.1 L of honey. How many servings can Alicia make if she has 0.8 L of honey? Use s for the number of servings. Choose all answers that are correct. A. 0.1s=0.8 B. 0.8/0.1=s C. 0.8 x s=0.1 D. s/0.8=0.1

OpenStudy (danjs):



OpenStudy (danjs):

\[\frac{ 0.8~L ~Honey }{ 1 }*\frac{ 1~serving }{ 0.1~L~Honey } = ~~servings\]

OpenStudy (danjs):

If you multiply both sides of that by 0.1 0.8 = 0.1*s

OpenStudy (danjs):

2 answers

OpenStudy (theedwardsfamily):

So its a and c?

OpenStudy (theedwardsfamily):

@DanJS

OpenStudy (michele_laino):

if we have 1 s for 0.1 L then we have 2 s for 0.2 L 3 s for 0.3 L 4 s for 0.4 L 5 s for 0.5 L 6 s for 0.6 L and so on, so?

OpenStudy (theedwardsfamily):

so the answers are a and c right @Michele_Laino ?

OpenStudy (michele_laino):

note that we have s=8 for 0.8 L then A and B, because you have to get s=8

OpenStudy (danjs):

A and B

OpenStudy (theedwardsfamily):

oh right got it! thanks

OpenStudy (danjs):

i wrote the answers above